OpenStudy (anonymous):

Find the period of the function. y = 5 cos one divided by twox

OpenStudy (anonymous):

Period (T) is inverse of frequency (f) Normal form of a sine/cosine is sine(2*pi*f*t), t=x (in your case) If you have cos(1/2x) match the arguments of the functions and solve 2*pi*f*x=1/2x

OpenStudy (anonymous):

i dont understand what your saying

OpenStudy (anonymous):

Sine and cosine are periodic functions When you study periodic functions they are always in the form of cos(2*pi*f*x) The function you gave is also in that form except the terms "2*pi" and "f" are already multiplied so you don't see it that easy By rule they are implicitly there So you can use "magic" (not really) of matching both arguments (whatever is inside the trig function) and you can get the frequency When you get the frequency (letter "f" ) you can get the period (is represented as "T") T = 1/f

OpenStudy (anonymous):

4π 5 pi divided by two five pi divided by two these are my answer choices

OpenStudy (anonymous):

\[2\pi*fx=(1/2)x\] \[2\pi*f=(1/2)\] \[f=1/(4\pi)\] \[T=1/f=1/(1/4\pi)=4\pi\]